# RDA-PSO: A computational method to quantify the diffusive dispersal of insects

### v1

@@ D.4 Application of the MDT model
Figure 5: Left panel: Error landscape produced by the grid search for the synthetic data of Table 4. The error $E$ is calculated over a uniform mesh of size $51\times 21$, with values of k in the interval $[50,150]$ and values of q in the interval $[10,50]$. Note that the surface is plotted using cubic interpolation over the actual data. Right panel: Contour plot of $E$ showing the estimated minimizer (red star) near k = 100 and q = 30. Here, the parameter p is set at ${\tt p}=0$ and the trap efficiency at ${\tt s}_{\tt e}=100$
@@ D.5.2 Particle swarm optimization
Once the general location of the error minimizer is identified by the systematic grid search procedure, the particle swarm optimization (PSO) method is applied to pinpoint an estimate of the location of the minimizer. Here, the PSO routine is run with 15 particles over 5 generations, in the parameter space region defined by k $\in[70,130]$ and q $\in[16,44]$. Out of 100 PSO runs, we keep the results whose error is less than the mean error plus one standard deviation, thus accepting 86% of the results. The average value of k = 99.85 with a standard deviation of 0.76 and the average value of q = 29.92 with a standard deviation of 0.21. The error ranges between $E=0.0856$ and $E=0.5562$. These numerically inferred average parameters are within a few percents of the actual values used to create the data (k = 100 and q = 30). If the direct problem in the RDA-PSO method is run with N=10,000 walkers instead of 100,000, we obtain k = $99.89\pm 0.996$ and q = $29.92\pm 0.29$ with the error ranging between 0.2464 and 0.6648.
@@ D.5.3 Estimation of the diffusion coefficient
Using the average value of k found by the RDA-PSO model leads to an estimate of the diffusion coefficient $D_{RDA}=299.1$ m2 per day, which is extremely accurate and presents a significant improvement on the $D_{TC}$ and $D_{ATC}$ estimates mentioned above. When implemented on synthetic data, the RDA-PSO method is thus able to estimate the diffusion coefficient with high accuracy.
@@ Appendix E Hainan study
The map of the village provided in [76] was used to plot Figure 6, in which each dot represents a house. Six concentric zones, centered either on the central (left panel of 6) or peripheral (right panel of Figure 6) release point are also shown. Recaptured numbers and recapture ratios of female mosquitoes released from both locations are listed in Tables 1 and 2 of the main text. The size of the zones associated with the release in the center of the village was adjusted to take into account that the house furthest from the release point did not catch any mosquitoes. Because the surface area of each zone significantly affects estimates of $D_{ATC}$, it is indeed important to only include capture sites that record mosquito presence. In addition, since this house is more than 15 meters away from the previous zone boundary, including it would lead to a zone of very large surface area with few collection sites, thereby leading to large and unrealistic corrections to the spatial (and temporal) ratios. Consequently, for the edge release, we also assume that this house did not catch any mosquitoes, i.e. that the 4 houses in the first 17 meters of zone 6 capture all of the mosquitoes for that zone. The furthest house is included in the RDA-PSO simulations since the surface area of the zones is not relevant for that method.
@@ E.1 Grid search
The first step of the optimization procedure consists in running simulations over a uniformly distributed sample of k and q while setting p = 0 and ${\tt s}_{\tt e}=100\%$. We take (k, q) in the range $[5,105]\times[3,43]$ for the center release and $[15,115]\times[3,43]$ for the edge release, with a total of 231 parameter pairs for each (21 values for k and 11 values for q). The error landscape, using equation (14), is shown in the left panel of Figure 7 for the center release and the right panel of Figure 7 for the edge release. This helps narrow down the PSO search to $[15,55]\times[5,30]$ and $[75,95]\times[5,40]$ respectively, with the ranges of p calculated so that no traps overlap and se ranging from 0.1% to 100%.
Figure 7: Error landscape calculated over a uniform mesh in the $({\tt k},{\tt q})$ space. Left panel: Error obtained for (k, q) in $[5,105]\times[3,43]$ where mosquitoes were released from the center. Right panel: Error obtained for (k, q) in $[15,115]\times[3,43]$ where mosquitoes were released from the edge. Note that the surface is plotted using cubic interpolation over the actual data
@@ Appendix F Cairns study
The zones associated to the experiment carried out in Cairns are drawn in figure 1 of [59], but their boundaries are not exactly circular, probably due to an anisotropic scaling of the figure during printing. In order to keep the correct number of capture sites in each zone, the $y$-scale of the map provided in [59] was slightly compressed on each side of the release site. This made the zones circular, while keeping each capture site in its associated zone. The result is shown in Figure 9; there are 7, 13, 19, and 27 capture sites in each zone respectively.
@@ F.2 Grid search and PSO results
For the grid search, we run forward simulations over a uniform sampling of k and q while setting p = 0 and ${\tt s}_{\tt e}=100\%$. We take (k, q) in the range $[20,320]\times[5,85]$, with a total of 527 parameter pairs for each (31 values for k and 17 values for q). The resulting error landscape, with the error defined in Equation (14), is displayed on the left panel of Figure 10. It shows a clear minimum, around which the initial conditions for the PSO step of the method are distributed.
@@
Figure 10: Left panel: Error landscape calculated over a uniform mesh in the $[20,320]\times[5,85]$ region of the (k, q) parameter space. The surface is plotted using cubic interpolation over the actual data. Note the larger variability of $E$ for very small values of q. Right panel: PSO results from 500 runs (as usual, outliers associated with a large error are removed; see main text for details). The optimal trap efficiency se is plotted as a function of the optimal inner radius q $\cdot$ h (blue circles, left panel) and of the optimal outer radius (q+p) $\cdot$ h (black circles, right panel) of each capture site
Appendix G Properties of TC, ATC, and MDT estimates
This section uses synthetic data to explore how the values of $D_{TC}$, $D_{ATC}$, and $D_{MDT}$ are affected by the MRR experimental setup.
@@ G.1 Uniform trap density: Sensitivity of TC estimates to experimental conditions
Figure 11: Estimates of $D_{TC}$ over 100 randomized trap locations. For each set of locations, the large blue dot is the average value and the crosses mark the maximum and minimum estimated values of $D_{TC}$. The $x$-axis shows the number of traps in the first zone, $nT_{1}$; the number of traps in the other zones is an odd multiple of $nT_{1}$, per the definition of $nT_{j}$. Parameters are k = 35, q = 20, p = 0, and $N=10,000$. Trap efficiency is 3%. The horizontal line represents the exact value of the diffusion coefficient $D$. Left panel: Temporal recapture ratios (see Equation (4) of the main text) are calculated daily. Right panel: Temporal recapture ratios are calculated hourly
@@ G.2 Heterogeneous trap density: Limitations of ATC and MDT estimates
Figure 13: Diffusion coefficient $D_{TC}$ estimates for 6 randomized full set of traps (blue dots) and 50 randomized subsets of these full sets (dark red stars). Estimates of $D_{ATC}$ with corrected spatial ratios (purple squares) and both spatial and temporal ratios (purple diamonds) are also shown. The green triangles correspond to $D_{MDT}$ values, with upper, left, and right triangles using uncorrected, corrected spatial, and corrected spatial and temporal ratios, respectively. In each case, the crosses indicate the minimum and the maximum estimates. The parameters are set at k = 35, q = 20, p = 0, and ${\tt s}_{\tt e}=3$, with $N=10,000$ walkers over 6 zones with $r_{zone}=15$ meters. The horizontal black line indicates the exact value of $D$ obtained from the fixed k parameter
When trap density is not uniform across capture zones, the time-corrected model needs to be adjusted with the correction factor of Equation (10) [38] or alternatively (18), thereby leading to the area-and-time-corrected model. To study the relation between the two models, we create 2 virtual MRR experiments with uniform trap distribution (each over 6 zones with $r_{zone}=15$ meters), the first one with $nT_{1}=4$ and the second with $nT_{1}=5$. We refer to each set of traps as a “full set” and randomly create 3 such full sets, for both $nT_{1}=4$ and $nT_{1}=5.$ Then we randomly choose 50 subsets of each full set to match the trap count of the Hainan center experiment, which has 4, 7, 6, 2, 0, and 1 traps in zones 1 through 6 respectively. We run simulations over each subset with $N=10,000$ walkers released from the center and parameters set at k = 35, q = 20, p = 0, and ${\tt s}_{\tt e}=3$. Each simulation is run over a period of 10 days and hourly time recapture ratios are collected. Figure 13 shows 3 sets of data for each virtual study, which include: the average of 10 estimates of $D_{TC}$ over the full set of traps (blue dots), the average of $D_{TC}$ estimates over the 50 subsets (dark red stars), the average of $D_{ATC}$ estimates over the 50 subsets when temporal ratios are not corrected (purple squares), and the average of $D_{ATC}$ estimates over the 50 subsets with corrected temporal ratios (purple diamonds), as well as estimates of $D_{MDT}$ using uncorrected ratios (green up triangles), corrected spatial ratios (green left triangles), and corrected spatial and temporal ratios (green right triangles). In each case, we also indicate the minimum and the maximum of the corresponding estimates of $D$.

### v2

@@ D.3 Application of the MDT model
D.4 Application of the time- and area-and-time-corrected models
Figure 5: Left panel: Error landscape produced by the grid search for the synthetic data of Table 7. The error $E$ is calculated over a uniform mesh of size $51\times 21$, with values of k in the interval $[50,150]$ and values of q in the interval $[10,50]$. Note that the surface is plotted using cubic interpolation over the actual data. Right panel: Contour plot of $E$ showing the estimated minimizer (red star) near k = 100 and q = 30. Here, the parameter p is set at ${\tt p}=0$ and the trap efficiency at ${\tt s}_{\tt e}=100$ %
@@ D.5.2 Particle swarm optimization
Once the general location of the error minimizer is identified by the systematic grid search procedure, the particle swarm optimization (PSO) method is applied to pinpoint an estimate of the location of the minimizer. Here, the PSO routine is run with 15 particles over 5 generations, in the parameter space region defined by k $\in[70,130]$ and q $\in[16,44]$. Out of 100 PSO runs, we keep the results whose error is less than the mean error plus one standard deviation, thus accepting 86% of the results. The average value of k = 99.85 with a standard deviation of 0.76 and the average value of q = 29.92 with a standard deviation of 0.21. The error ranges between $E=0.0856$ and $E=0.5562$ (after removing the outliers). These numerically inferred average parameters are within a few percents of the actual values used to create the data (k = 100 and q = 30). If the direct problem in the RDA-PSO method is run with N=10,000 walkers instead of 100,000, we obtain k = $99.89\pm 0.996$ and q = $29.92\pm 0.29$ with the error ranging between 0.2464 and 0.6648.
@@ D.5.3 Estimation of the diffusion coefficient
Using the average value of k found by the RDA-PSO model, k = 99.85, leads to an estimate of the diffusion coefficient $D_{RDA}=299.1$ m2 per day, which is extremely accurate and presents a significant improvement on the $D_{MDT}$, $D_{TC}$, and $D_{ATC}$ estimates mentioned above. When implemented on synthetic data, the RDA-PSO method is thus able to estimate the diffusion coefficient with high accuracy, as shown in Table 8, which summarizes all the diffusion coefficient estimates obtained in this section, along with their signed relative error.
[…]
estimate relative error
[…]
$D_{MDT}$ 278.3 $7.23\%$
$D_{MDT}$ (corrected) 320.3 $6.8\%$
$D_{TC}$ 234.6 $21.8\%$
$D_{ATC}$ 331.4 $10.47\%$
$D_{RDA}$ 299.1 $0.3\%$
[…]
@@ Appendix E Hainan study
The map of the village provided in [80] was used to plot Figure 6, in which each dot represents a house. Six concentric zones, centered either on the central (left panel of 6) or peripheral (right panel of Figure 6) release point are also shown. Recaptured numbers and recapture ratios of female mosquitoes released from both locations are listed in Tables 3 and 4 of the main text. The size of the zones associated with the release in the center of the village was adjusted to take into account that the house furthest from the release point did not catch any mosquitoes. Because the surface area of each zone significantly affects estimates of $D_{ATC}$, it is indeed important to only include capture sites that record mosquito presence. In addition, since this house is more than 15 meters away from the previous zone boundary, including it would lead to a zone of very large surface area with few capture sites, thereby leading to large and unrealistic corrections to the spatial (and temporal) ratios. Consequently, for the edge release, we also assume that this house did not catch any mosquitoes, i.e. that the 4 houses in the first 17 meters of zone 6 capture all of the mosquitoes for that zone. The furthest house is included in the RDA-PSO simulations since the surface area of the zones is not relevant for that method.
@@ E.1 Grid search
The first step of the optimization procedure consists in running simulations over a uniformly distributed sample of k and q while setting p = 0 and ${\tt s}_{\tt e}=100\%$. We take (k, q) in the range $5,105]\times[3,43]$ for the center release and $15,115]\times[3,43]$ for the edge release, with a total of 231 parameter pairs for each (21 values for k and 11 values for q). The error landscape, using equation (14), is shown in the left panel of Figure 7 for the center release and the right panel of Figure 7 for the edge release. This helps narrow down the PSO search to $15,55]\times[5,30]$ and $75,95]\times[5,40]$ respectively, with the ranges of p calculated so that no traps overlap and se ranging from 0.1% to 100%.
Figure 7: Error landscape calculated over a uniform mesh in the $({\tt k},{\tt q})$ space. Left panel: Error obtained for (k, q) in $5,105]\times[3,43]$ where mosquitoes were released from the center. Right panel: Error obtained for (k, q) in $15,115]\times[3,43]$ where mosquitoes were released from the edge. Note that the surface is plotted using cubic interpolation over the actual data
@@ Appendix F Cairns study
The zones associated to the experiment carried out in Cairns are drawn in figure 1 of [63], but their boundaries are not exactly circular, probably due to an anisotropic scaling of the figure during printing. In order to keep the correct number of capture sites in each zone, the $y$-scale of the map provided in [63] was slightly compressed on each side of the release site. This made the zones circular, while keeping each capture site in its associated zone. The result is shown in Figure 9; there are 7, 13, 19, and 27 capture sites in each zone respectively. The study site includes streets, which could limit mosquito spread or serve as corridors that facilitate dispersal. Because our simulations assume a homogeneous landscape away from capture sites, the estimates of $D$ we obtain may not be accurate. They can however be compared to one another, which is the purpose of the present discussion.
@@ F.2 Grid search and PSO results
For the grid search, we run forward simulations over a uniform sampling of k and q while setting p = 0 and ${\tt s}_{\tt e}=100\%$. We take (k, q) in the range $20,320]\times[5,85]$, with a total of 527 parameter pairs for each (31 values for k and 17 values for q). The resulting error landscape, with the error defined in Equation (14), is displayed on the left panel of Figure 10. It shows a clear minimum, around which the initial conditions for the PSO step of the method are distributed.
@@
Figure 10: Left panel: Error landscape calculated over a uniform mesh in the $20,320]\times[5,85]$ region of the (k, q) parameter space. The surface is plotted using cubic interpolation over the actual data. Note the larger variability of $E$ for very small values of q. Right panel: PSO results from 500 runs (as usual, outliers associated with a large error are removed; see main text for details). The optimal trap efficiency se is plotted as a function of the optimal inner radius q $\cdot$ h (blue circles, left panel) and of the optimal outer radius (q+p) $\cdot$ h (black circles, right panel) of each capture site
Appendix G Properties of MDT, TC, and ATC estimates
This section uses synthetic data to explore how the values of $D_{MDT}$, $D_{TC}$, and $D_{ATC}$ are affected by the MRR experimental setup. A comparison with the results of the RDA-PSO method (based on 20 independent runs) is also provided.
@@ G.1 Uniform trap density: Sensitivity of TC estimates to experimental conditions
Figure 11: Estimates of $D_{TC}$ for different trap densities. The $x$-axis shows the number of traps in the first zone, $nT_{1}$; the number of traps in the other zones is an odd multiple of $nT_{1}$, per the definition of $nT_{j}$. For each $nT_{1}$ value, there are 100 sets of randomized trap locations. The large blue dot is the average value and the crosses mark the maximum and minimum estimated values of $D_{TC}$. Parameters are k = 35, q = 20, p = 0, and $N=10,000$. Trap efficiency is 3%. The horizontal line represents the exact value of the diffusion coefficient $D$. Left panel: Temporal recapture ratios (see Equation (4) of the main text) are calculated daily. Right panel: Temporal recapture ratios are calculated hourly
@@ G.2 Heterogeneous trap density: Limitations of ATC and MDT estimates
Figure 13: Diffusion coefficient $D_{TC}$ estimates for 6 randomized full set of traps (blue dots) and 50 randomized subsets of these full sets (dark red stars). Estimates of $D_{ATC}$ with corrected spatial ratios (purple squares) and both spatial and temporal ratios (purple diamonds) are also shown. The green triangles correspond to $D_{MDT}$ values, with upper, left, and right triangles using uncorrected, corrected spatial, and corrected spatial and temporal ratios, respectively. In each case, the crosses indicate the minimum and the maximum estimates. The parameters are set at k = 35, q = 20, p = 0, and ${\tt s}_{\tt e}=3$ %, with $N=10,000$ walkers over 6 zones with $r_{zone}=15$ meters. The horizontal black line indicates the exact value of $D$ obtained from the fixed k parameter
When trap density is not uniform across capture zones, the time-corrected model needs to be adjusted with the correction factor of Equation (12) [41] or alternatively (18), thereby leading to the area-and-time-corrected model. To study the relation between the two models, we create 2 virtual MRR experiments with uniform trap distribution (each over 6 zones with $r_{zone}=15$ meters), the first one with $nT_{1}=4$ and the second with $nT_{1}=5$. We refer to each set of traps as a “full set” and randomly create 3 such full sets, for both $nT_{1}=4$ and $nT_{1}=5.$ Then we randomly choose 50 subsets of each full set to match the trap count of the Hainan center experiment, which has 4, 7, 6, 2, 0, and 1 traps in zones 1 through 6 respectively. We run simulations over each subset with $N=10,000$ walkers released from the center and parameters set at k = 35, q = 20, p = 0, and ${\tt s}_{\tt e}=3$ %. Each simulation is run over a period of 10 days and hourly time recapture ratios are collected. Figure 13 shows 3 sets of data for each synthetic study, which include: the average of 10 estimates of $D_{TC}$ over the full set of traps (blue dots), the average of $D_{TC}$ estimates over the 50 subsets (dark red stars), the average of $D_{ATC}$ estimates over the 50 subsets when temporal ratios are not corrected (purple squares), and the average of $D_{ATC}$ estimates over the 50 subsets with corrected temporal ratios (purple diamonds), as well as estimates of $D_{MDT}$ using uncorrected ratios (green up triangles), corrected spatial ratios (green left triangles), and corrected spatial and temporal ratios (green right triangles). In each case, we also indicate the minimum and the maximum of the corresponding estimates of $D$.
@@
G.3 Comparison with RDA-PSO estimates
This section compares estimates returned by the MDT, TC, ATC, and RDA-PSO models for synthetic experiments with trap configurations similar to those of Section G.2. We randomly generate 10 subsets of trap locations from the full set used for Simulation 1 of the previous section, for which $nT_{1}=4$. Input data consist of the recapture ratios of $N=3,000$ walkers, with parameters set at k = 35, q = 20, p = 0, and ${\tt s}_{\tt e}=3$ %, over 6 zones with $r_{zone}=15$ meters. For each subset, we calculate $D_{MDT}$, $D_{TC}$, $D_{ATC}$, using both the spatially and temporally corrected ratios (for the MDT and ATC estimates). To calculate $D_{RDA}$, we obtain optimal values of k from 20 PSO runs with $N_{p}=15$ particles and $N_{g}=5$ generations since this is a 2-parameter optimization, remove outliers, find the average value of the remaining k’s, and substitute the result in equation (13). Figure 14 compares the resulting estimates with the known value of $D$, for each of the 10 subsets of trap configurations. Minimum and maximum values of the signed relative errors for each of the 4 methods are listed in Table 10. These results show that, in situations where the true value of the diffusion coefficient is known, the RDA-PSO outperforms the other three methods. In addition, the RDA-PSO estimate has a relative error of less than 11% when 20 PSO simulations are used.
Figure 14: Comparison of the diffusion coefficient estimates for the four approaches introduced in this article, for 10 different subsets of randomly generated trap locations. The estimates for $D_{MDT}$ and $D_{ATC}$ are computed with both spatial and temporal corrections. The $D_{RDA}$ estimate is calculated from 20 PSO runs after removing outliers. The parameters are set at k = 35, q = 20, p = 0, and ${\tt s}_{\tt e}=3$ %, with $N=3,000$ walkers over 6 zones with $r_{zone}=15$ meters. The horizontal black line indicates the exact value of $D$ obtained from the fixed k parameter
[…]
minimum relative error maximum relative error
(from the exact $D$) (from the exact $D$)
$D_{MDT}$ (sp.+temp.) $29.64\%$ $20.6\%$
$D_{TC}$ $49.7\%$ $41.69\%$
$D_{ATC}$ (sp.+temp.) $19.35\%$ $9.21\%$
$D_{RDA}$ $10.26\%$ $6.05\%$
[…]
Appendix H Effect of trap density on recapture ratios
Recapture ratios, $\tau_{i}$ and $\sigma_{j}$, are sometimes affected by the value of trap efficiency, se, depending on how many traps are used in an experiment. If the trap density is low, recapture ratios remain robust to changes in se. However, if the trap density is high, recapture ratios change depending on the value of se. This is illustrated in the following virtual experiment. We consider 6 concentric zones (separated by $r_{zone}=15$ meters, as in Appendix G) and a uniform trap density. In the first scenario, we place 1 trap in the first zone, that is, $nT_{1}=1$; whereas in the second scenario, $nT_{1}=4$. We release $N=10,000$ mosquitoes and for every forward run, we vary the trap efficiency to see how the recapture percentages change. The left panel of Figure 15 shows that when less traps are used ($nT_{1}=1$), the percentages are close to each other, no matter the value of se. On the other hand, the right panel of Figure 15 shows that the recapture percentages vary considerably with se when the number of traps is high. Since the traps are uniformly distributed, having 4 traps in zone 1 means there are 144 traps in the entire field of study. In that case, the right panels of Figure 15 indicate that most of the walkers are captured on the first two days and in the first two zones when se is high, suggesting the associated MRR experiment is badly designed. In other words, more is not better in this case, and the number of traps used for a MRR experiment should be chosen so that the spatial and temporal ratios remain non-negligible in most of the zones, for the entire duration of the experiment.
Figure 15: Recapture ratios, $\sigma_{j}$ (top plot) and $\tau_{i}$ (bottom plot), corresponding to virtual experiments with different trap efficiency, se. Traps are distributed uniformly according to zone areas, with 6 concentric zones 1515 apart. Left panel: The first zone has 1 trap (low trap density). Right panel: The first zone has 4 traps (high trap density)